Waterbury Insurance Company wants to study the relationship between the amount of fire damage and the distance between the bua-2. Is there a direct or indirect relationship between the distance from the fire station and the amount of fire damage? TheC-1. Determine and interpret the coefficient of determination. (Round your answer to 3 decimal places.) Coefficient of determd-1. Determine the correlation coefficient. (Round your answer to 3 decimal places.) Correlation coefficient d-2. Choose thed-3. How did you determine the sign of the correlation coefficient? It is because the slope is e-1. State the decision rule fe-2. Compute the value of the test statistic. (Round your answer to 2 decimal places.) Value of the test statistic e-3. Is th
Answer #1

a-1)          
y=1381345+300415*x          
          
a-2)          
direct          
          
b) estimated damage=1381345+300415*7=   3484250      
c-1) coefficient of determination = SSR/SST=1805.5782/3081.0716=   0.586      
c-2) 58.6%          
d-1) correlation coefficient,r= √0.586=   0.766      
d-2) positive          
d3) positive, positive          
e1) reject Ho if t <   -2.763   or t >    2.763
e-2) value of test stat= 6.3          
e-3) reject, significant